Apples are collected in a basket for 6 people. one third, one forth, one eighth, and one fifth of the apples are given to four people, respectively. The fifth person gets 10 apples and one apple remains for the sixth person. find an original number of apples in the basket???.
Answer by scott8148(6628)

(x/3) + (x/4) + (x/5) + (x/8) = x - 11

multiplying by 120 (LCD) ___ 40x + 30x + 24x + 15x = 120x - 1320 ___ 109x = 120x = 1320

-11x = -1320 ___ x = 120